# Partial Knowledge Distillation for Alleviating the Inherent Inter-Class Discrepancy in Federated Learning

### v1

@@ III-A Overview
In this work, we introduce a federated partial knowledge distillation (PKD) method to mitigate the impact of inherent ICD and enhance the classification accuracy of weak classes. The flow of the proposed method is outlined in Algorithm 1. The initial model first undergoes several rounds of federated learning (e.g. 10 to 20 rounds for FashionMNIST). After the warmup stage, it enters an expert learning phase. Weak classes are identified at the edge by examining the output probability distribution obtained from a feedforward pass on the local training set. Fig. 6 shows the predicted probabilities for classes other than the true class. Each curve represents the probability of samples from one class being misclassified as another class. As depicted in the figure, after the short warmup phase, the chance of misclassifying one weak class as another within the same group—such as a ‘T-shirt’ being predicted as a ‘shirt’ or vice versa—is significantly higher compared to other cases. This distinction can help to identify the weak class groups. The identified groups of weak classes demonstrate a notable consistency across different clients. $g_{i}$ in Line 7 refers to a group of classes that can easily be confused with each other. For example, in FashionMNIST, $g_{1}$ consists of {‘T-shirt’, ‘shirt’}, and $g_{2}$ comprises of {‘pullover’, ‘coat’, and ‘shirt’}. In the case of CIFAR-10, there is only one group including {‘bird’, ‘cat’, ‘deer’ and ‘dog’}. Once a group of weak classes is identified ($g_{i}$), a specialized expert model is trained on a subset of the training data ($D_{g_{i}}$) to extract and discern the subtle feature differences among them. All experts are trained within the FL framework, ensuring that no data is uploaded to the cloud. This process takes 25 local training rounds for the FashionMNIST dataset. Typically, one or two experts are adequate to address all the weak classes as described in Section II-B.
@@
Figure 6: The Probability of Samples from One Class Being Misclassified as Another Class. (FashionMNIST: 90 Curves in Total)
III-B Partial Knowledge Distillation
The complete pseudo-code of the partial KD process is given in Algorithm 2. The overall framework is similar to the baseline approach of FedAvg [1], with the exception of the loss calculation described within the dashed box. At each round, the global model is broadcast to $K$ randomly selected clients. Each client trains the model locally for $E$ epochs using its own dataset (Line 5). Then, the local models are sent back and aggregated on the server (Line 7) as described in [1].
During the local training phase on client $k$, the prediction results are analyzed to detect any instances of misclassification among certain classes. An illustration example using the FashionMNIST dataset is presented in Fig. 7. Once misclassification within group $g_{i}$ occurs, the corresponding expert model $exp_{i}$ is triggered. A KD loss is then computed (Line 17 and 18) using the KL divergence [22] between $p_{s}$ and $p_{e}$:
$L_{KD}=D_{KL}(p_{s}\parallel p_{e})$ (1)
where $p_{s}$ and $p_{e}$ represent the predicted probabilities obtained from the student model and the expert model, respectively. The probabilities are calculated using the Softmax function with a temperature parameter $T$, formulated as follows:
$p_{i}=\frac{e^{z_{i}/T}}{\sum_{j}e^{z_{j}/T}}$ (2)
where $z_{i}$ denotes the output logit. Since the partial KD method targets specific classes, $p_{s}$ and $p_{e}$ only involve a subset of classes, as illustrated in Fig. 7. In the case where no expert model is activated ($exp_{i}$ = None in Line 16), $L_{KD}$ is simply set to zero. The complete loss function is formulated as Line 21, where $L_{CE}$ denotes the standard cross-entropy loss, and $\lambda$ is a hyperparameter.
III-C Overhead Analysis
@@
$U_{T_{2}}=E_{exp}\cdot\sum_{i}n_{g_{i}}\cdot(U_{FP}+U_{BP})=\frac{E_{exp}}{E}\cdot\frac{\sum_{i}n_{g_{i}}}{N}\cdot U$ (5)
@@
$U_{T_{3}}=U+n_{KD}^{t}\cdot U_{FP}\approx(1+\frac{1}{3}\cdot\frac{n_{KD}^{t}}{E\cdot N})\cdot U$ (6)
@@
Figure 7: An Illustration Example for the Partial KD Process.
IV Experimental Results
IV-A Experimental Setup
We evaluate the performance of our approach on three commonly used datasets in the field of FL: MNIST, FashionMNIST, and CIFAR-10. The data are partitioned in the same way as described in Section II-B. Two widely used neural networks, LeNet-5 and ConvNet-S as in [1], are deployed for the MNIST and FashionMNIST datasets, respectively. For the CIFAR-10 dataset, we employ larger networks including VGG-9 [23] and ConvNet-L (similar to ConvNet-S but with one more 5$\times$5 Convolutional layer). In this section, we compare the performance of the PKD method with the baseline benchmark FedAvg, as well as several prior techniques designed to mitigate the ICD problem resulting from class imbalance. These techniques include focal loss [11] and data augmentation [20, 14]. Stochastic Gradient Descent is employed for optimization, with a learning rate set to 0.01. The number of local epochs and the batch size is optimized to 5 and 50, respectively. The models are trained for 100 rounds on MINIST and 300 rounds on the other datasets. The experts’ network structure is identical to the global model’s. The number of experts aligns with the weak class subsets detailed in Section II-B: two for MNIST and FashionMNIST, and one for CIFAR-10. During the PKD process, the hyperparameter $T$ is set to 3. $\lambda$ is tailored to 1 for both MNIST and FashionMNIST, and to 0.5 for CIFAR-10.
IV-B Performance of the Proposed Method
@@
Table I presents the performance of our proposed method under conditions where the class distribution is balanced both globally and locally. In the baseline scenario, the inter-class accuracy discrepancy is 1.42%, 45.1% and 27.1% for MNIST, FashionMNIST and CIFAR-10, respectively. The proposed PKD method can improve the accuracy of the worst-performing classes by 0.4%, 10.7% and 6.6%, respectively, without comprising the global average accuracy. Furthermore, these improvements are more pronounced when the local class distribution is extremely unbalanced, as shown in Table I. In such cases, the minimum class-wise accuracy increases by 60.5% and 51.8% on MNIST and FashionMNIST, respectively. As shown in Fig. 8, the accuracy of weak classes becomes significantly more stable compared to the baseline case.
@@
Figure 8: Accuracy of the Worst-Performing Class ‘Shirt’ During Training. (FashionMNIST, Local Class Imbalance as in Fig. 5)
IV-C Comparison with Existing Techniques
@@
Table II compares our proposed method with several existing techniques that address the ICD problem brought by class imbalance. An effective approach to address the class imbalance issue is to offset the negative impact brought by the uneven class distribution [12, 11]. It can be done by either re-sampling the minority classes or assigning larger weights to their loss values while calculating the empirical risk. Here we adopt a similar idea and test its performance for addressing the inherent ICD problem. Instead of up-weighting the minority classes, we dynamically enhance the weights of the inherent weak classes when calculating the loss value. The results indicate that the minimum class-wise accuracy may not be improved by this method. In [20] and [14], data augmentation is applied at the sample level and feature level, respectively, to implicitly enhance data diversity. These two approaches are commonly utilized to improve the model’s performance on minority classes. The results suggest that feature-level data augmentation can modestly mitigate the inherent ICD issue, leading to a 2.30% improvement in the accuracy of the worst-performing class.
IV-D Overhead Analysis

### v2

@@ 3.1. Overview
In this work, we introduce a federated partial knowledge distillation (PKD) method to mitigate the impact of inherent ICD and enhance the classification accuracy of weak classes. The flow of the proposed method is outlined in Algorithm 1. The initial model first undergoes several rounds (e.g., 10 to 20) of federated learning. After the warmup stage, it enters an expert learning phase (Line 6). Weak classes are first identified at the edge by examining the output probability distribution obtained from a feedforward pass on the local training set. For illustrative purposes, we use the FashionMNIST dataset as an example. Fig. 7(a) shows the predicted probabilities for classes other than the true class. Each curve represents the probability of samples from one class being misclassified as another class. As depicted in the figure, after the short warmup phase, the chance of misclassifying one weak class as another within the same group—such as a ‘T-shirt’ being predicted as a ‘shirt’ or vice versa—is significantly higher compared to other cases. This distinction can help to identify the weak class groups. Specifically, the eight highlighted curves in Fig. 7(a) correspond to four pairs of weak classes, namely {‘shirt’, ‘T-shirt’}, {‘shirt’, ‘coat’}, {‘shirt’, ‘pullover’}, and {‘coat’, ‘pullover’}. These four weak class pairs form two groups, as shown in the figure. Fig. 7(b) further presents the similarities in high-level features ($f$) between each pair of classes at training round 10$\sim$20. It is worth noting that both Fig. 7(a) and (b) reveal pronounced disparities between weak classes and others, even during the early training phase.
Figure 7. (a) The Probability of Samples from One Class Being Misclassified as Another (FashionMNIST, 10$\times$9 Curves in Total).    (b) Euclidean Distance between High-Level Features ($f$) of each Class Pair at Round 10$\sim$20.
The identified weak class groups are denoted as $g_{i}$ in Line 7. For example, in FashionMNIST, $g_{1}$ consists of {‘T-shirt’, ‘shirt’}, and $g_{2}$ comprises of {‘pullover’, ‘coat’, and ‘shirt’}. In the case of CIFAR-10, there is only one weak class group {‘bird’, ‘cat’, ‘deer’ and ‘dog’}. Once a group of weak classes is identified ($g_{i}$), a specialized expert model is trained on a subset of the training data ($D_{g_{i}}$) to extract and discern the subtle feature differences among them. ($D_{g_{i}}$ in Line 10 refers to the samples from class group $g_{i}$.) For simplicity, the expert model adopts the same network structure as the global model, with the exception of the last layer. The last layer has fewer neurons, as the expert is trained for specific classes. All experts are trained within the FL framework, ensuring that no data is uploaded to the cloud. This process takes 25 local training rounds for the FashionMNIST dataset. The overhead of expert training for various benchmarks is analyzed in Section 3.3 and Section 4.5. Typically, one or two experts are adequate to effectively improve the accuracy of the worst-performing classes and alleviate the inherent ICD problem.
@@
3.2. Partial Knowledge Distillation
The complete pseudo-code of the PKD process is given in Algorithm 2. The overall framework is similar to the standard baseline, FedAvg (McMahan et al., 2017), with the exception of the loss calculation described in Line 22. At each round, the global model is broadcast to $K$ randomly selected clients. Each client trains the model locally for $E$ epochs using its own dataset (Line 5). Then, the local models are sent back and aggregated on the server (Line 7) as described in (McMahan et al., 2017).
During the local training phase on client $k$, the prediction results ($preds$) are analyzed to detect any instances of misclassification among certain classes in Line 14. The indices of these misclassified samples are denoted as $sample\_id$. An illustration example is presented in Fig. 8. Once misclassification within group $g_{i}$ occurs, the corresponding expert model $exp_{i}$ is triggered. The experts’ prediction results on the misclassified samples ($x[{sample\_id}]$) are denoted as $p_{e}$ (Line 17). A PKD loss is then computed using the KL divergence (Hinton et al., 2015) between $p_{s}$ and $p_{e}$ in Line 19:
(1) $L_{PKD}=D_{KL}(p_{s}\parallel p_{e})$
where $p_{s}$ represents the corresponding predicted probabilities obtained from the student model, as illustrated in Fig. 8. The probabilities $p_{e}$ and $p_{s}$ are calculated using the Softmax function with a temperature parameter $T$, formulated as follows:
(2) $p_{i}=\frac{e^{z_{i}/T}}{\sum_{j}e^{z_{j}/T}}$
where $z_{i}$ denotes the output logit. Since the PKD method targets specific classes, $p_{s}$ and $p_{e}$ only involve a subset of classes, e.g., 3 classes in Fig. 8. In the case where no expert model is activated ($exp$ = None in Line 16), $L_{PKD}$ is simply set to zero. The complete loss function is formulated as Line 22 in Algorithm 2, where $L_{CE}$ denotes the standard cross-entropy loss, and $\lambda$ is a hyperparameter.
Figure 8. An Illustration Example for the Partial KD Process.
3.3. Overhead Analysis
@@
(5) $U_{T_{2}}=E_{exp}\cdot\sum_{i}n_{g_{i}}\cdot(U_{FP}+U_{BP})=\frac{E_{exp}}{E}\cdot\frac{\sum_{i}n_{g_{i}}}{N}\cdot U$
@@
(6) $U_{T_{3}}=U+n_{KD}^{t}\cdot U_{FP}\approx(1+\frac{1}{3}\cdot\frac{n_{KD}^{t}}{E\cdot N})\cdot U$
@@
4. Experimental Results
4.1. Experimental Setups
Datasets. We evaluate the proposed method on several widely-used datasets in FL, including the simple datasets MNIST and FashionMNIST, as well as the more complex datasets CIFAR-10 and CIFAR-100 (Krizhevsky, 2009). We also adopt the more realistic and naturally partitioned dataset FEMNIST (Caldas et al., 2018) for evaluation.
Data partitioning strategies. To make a comprehensive analysis, we employ four data partitioning strategies. 1) Local class-balanced partition: Each client’s dataset maintains strictly equal sample sizes across all classes, ensuring uniform local class distribution. 2) Pathological partition (McMahan et al., 2017): Each client contains data samples from only a small subset of classes (e.g., 1-2 classes), leading to severe distribution skew. Unless otherwise stated, we adopt a single-class-per-client assignment by default to ensure a clear and straightforward comparison of class-wise accuracy. 3) The Dirichlet partition (Yurochkin et al., 2019): Sample allocation follows a Dirichlet distribution $Dir(\alpha)$ (Minka, 2000) to generate a range of non-i.i.d. data scenarios, enabling comprehensive analysis of the inherent ICD phenomenon and evaluation of the PKD method under varying heterogeneity levels. 4) FEMNIST’s Native Partition: The dataset’s native writer-based partition is preserved to simulate practical FL scenarios. Global class balance across all clients is preserved in all experiments except for FEMNIST, which maintains its natural data distribution.
Models. Similar to (McMahan et al., 2017; Wang et al., 2020; Li et al., 2023), LeNet-5 is adopted for the MNIST, FashionMNIST, and FEMNIST datasets by default. We also use the same lightweight network ConvNet-S as in (McMahan et al., 2017) for comparison. (ConvNet-L is simply an extended version of ConvNet-S with an additional 5$\times$5 convolutional layer.) For CIFAR-10 and CIFAR-100, we use ResNet-56 (He et al., 2016) by default and also evaluate with the shallower network VGG-9 (Simonyan and Zisserman, 2014; Wang et al., 2020). As mentioned in Section 3.1, the expert model adopts the same network structure as the global model, with the exception of the last layer. The number of output neurons is equal to the size of the weak class subset. The number of experts aligns with the number of weak class subsets detailed in Section 2.2: two for MNIST and FashionMNIST, and one for CIFAR-10. Since CIFAR-100 comprises 100 classes, multiple weak class groups exist. We select the $G$ worst-performing groups based on the lowest average accuracy.
Table 1. Performance of PKD under the Local Class-Balanced Scenario
#Client Dataset MNIST FashionMNIST CIFAR-10 Method FedAvg PKD $\Delta$ FedAvg PKD $\Delta$ FedAvg PKD $\Delta$ 10 Max. 99.73 99.73 0.00 97.30 97.60 0.30 93.10 93.00 -0.10 Ave. 99.28 99.32 0.04 89.29 89.50 0.21 83.78 84.93 1.15 Min. 98.31 98.71 0.40 68.40 76.60 8.20 59.60 71.90 12.30 Worst “9” “9” “Shirt” “Shirt” “Cat” “Cat” 100 Max. 99.28 99.29 0.01 96.70 96.60 -0.10 80.90 82.70 1.80 Ave. 98.11 98.34 0.23 84.66 84.72 0.06 68.52 70.11 1.59 Min. 96.72 96.92 0.20 51.60 62.30 10.70 46.70 53.00 6.30 Worst “9” “9” “Shirt” “Shirt” “Cat” “Cat”
Compared Methods. In this section, we compare the performance of the PKD method with the standard baseline, FedAvg, as well as several prior techniques (Lin et al., 2020; Shuai et al., 2022; Devries and Taylor, 2017; Sahu et al., 2018; Ni et al., 2022) under various data partitioning setups. These advanced techniques include focal loss (Lin et al., 2020) and data augmentation (Shuai et al., 2022; Devries and Taylor, 2017), both designed to mitigate the ICD resulting from class imbalance. We also compare with two other methods, FedProx (Sahu et al., 2018) and FedCodl (Ni et al., 2022), under the Dirichlet partitions. These methods are designed to alleviate the data heterogeneity problem in FL by either constraining local updates or incorporating global knowledge into the local training process through KD. Our analysis reveals that while current approaches effectively alleviate general data heterogeneity, they remain insufficient in addressing the inherent ICD problem.
Hyperparameters. Stochastic Gradient Descent (SGD) is employed for optimization, with a learning rate of 0.01. Same as FedAvg (McMahan et al., 2017), the number of local epochs and the batch size are set to 5 and 50, respectively. The networks are trained for 100 rounds on MNIST and up to 300 rounds on other datasets. The total number of clients is set to 10 by default. We also scale this default value to 100 across various benchmarks and vary the fraction of clients selected at each round to evaluate PKD under different setups. Three fractions are used on the FEMNIST dataset, including 0.1, 0.5, and 1.0. During the PKD process, the hyperparameter $T$ is set to 5. The value of $\lambda$ is set to 1.0 for MNIST, FashionMNIST, and FEMNIST, and to 0.5 for CIFAR-10 and CIFAR-100.
Implementation. All experiments are implemented in Python using the PyTorch framework. Training is conducted on a server running Ubuntu, equipped with 4 NVIDIA RTX 3090 GPUs, 2 Intel Xeon Platinum 8269CY CPUs, and 128 GB of RAM.
4.2. Performance of the Proposed Method
Table 2. Results for Other Neural Networks under the Local Class-Balanced Scenario
Dataset FashionMNIST CIFAR-10 Network ConvNet-S VGG-9 Method FedAvg PKD $\Delta$ FedAvg PKD $\Delta$ Max. 98.20 98.20 0.00 92.50 92.60 0.10 Ave. 90.43 90.92 0.49 83.04 83.48 0.44 Min. 70.50 75.30 4.80 65.10 72.30 7.20 Worst “Shirt” “Shirt” “Cat” “Cat”
Table 1 presents the performance of our proposed method under local class-balanced scenario as described in Section 4.1. In the baseline scenario with 10 clients, the inter-class accuracy discrepancy is 1.42%, 28.9% and 33.5% for MNIST, FashionMNIST and CIFAR-10, respectively. The proposed PKD method can improve the accuracy of the worst-performing classes by 0.4%, 8.2% and 12.3%, respectively, without comprising the global average accuracy. The overall model performance degrades as the number of clients increases to 100. In this case, PKD can improve the worst class-wise accuracy by 0.2%, 10.7%, and 6.3% on MNIST, FashionMNIST, and CIFAR-10, respectively. It is worth noting that improving the accuracy of weak classes usually does not compromise the model performance on other classes. The accuracy of dominant classes remains largely unchanged or is even slightly improved. As a result, the global average accuracy also increases. As mentioned in Section 4.1, we employ LeNet-5 as the baseline architecture for evaluations on MNIST and FashionMNIST, while utilizing ResNet-56 for CIFAR-10. For a comprehensive comparison, Table 2 further presents the performance of PKD for additional network architectures.
Table 3. Performance of PKD under the Pathological Data Partition Scenario
Dataset MNIST FashionMNIST CIFAR-10 #Class/Client 1 1 2 Method FedAvg PKD $\Delta$ FedAvg PKD $\Delta$ FedAvg PKD $\Delta$ Max. 96.02 99.03 3.01 91.20 91.40 0.20 72.10 80.30 8.20 Ave. 75.85 84.45 8.60 47.39 71.03 23.64 38.64 41.42 2.78 Min. 7.90 68.38 60.48 0.00 51.80 51.80 2.90 16.10 13.20 Worst “8” “9” “Shirt” “Shirt” “Bird” “Cat”
@@ 4.2. Performance of the Proposed Method
Table  3 shows the performance of PKD under the pathological partition scenario. The number of classes per client (#Class/Client) is set to 1 for MNIST and FashionMNIST, and to 2 for CIFAR-10. Similar to the results reported in (Li et al., 2022), the global accuracy becomes low under such an extremely unbalanced distribution. Compared to local class-balanced scenarios, performance improvements for the worst-performing classes are more pronounced, as shown in Table  3. The minimum class-wise accuracy increases by 60.5% on MNIST, 51.8% on FashionMNIST, and 13.2% on CIFAR-10, respectively. As a result, the global average accuracy is also improved. As shown in Fig. 9, the accuracy of weak classes becomes significantly more stable compared to the baseline case.
Table  4 presents the experimental results under the Dirichlet partition scenario. The concentration parameter $\alpha$ in $Dir(\alpha)$ is set to 0.5. The overall model performance is better than in the pathological partition scenario. However, the baseline methods still exhibit 3.01% inherent ICD on MNIST, 43.8% on FashionMNIST, and 37.6% on CIFAR-10. Our PKD approach improves the minimum class-wise accuracy by 0.39%, 10.10%, and 6.10%, respectively, without affecting the average accuracy.
Table 4. Performance of PKD under the Dirichlet Partition Scenario ($Dir~{}(0.5)$)
Dataset MNIST FashionMNIST CIFAR-10 Method FedAvg PKD $\Delta$ FedAvg PKD $\Delta$ FedAvg PKD $\Delta$ Max. 99.64 99.64 0.00 96.90 96.90 0.00 92.90 93.80 0.90 Ave. 98.72 98.84 0.12 86.05 86.20 0.15 76.38 77.92 1.54 Min. 96.63 97.02 0.39 53.10 63.20 10.10 55.30 61.40 6.10 Worst “9” “9” “Shirt” “Shirt” “Dog” “Cat”
@@
Figure 9. Accuracy of the Worst-performing Class ‘Shirt’ during Training. (FashionMNIST, #Class/Client=1)
Table  5 presents the results on the FEMNIST dataset based on its native partition. The dataset comprises of over 3000 users and 62 classes, including digits, uppercase, and lowercase letters. To make a direct comparison with the MNIST dataset, we retain only the digit samples and their corresponding users. In realistic FL scenarios, not all users participate in every communication round. To simulate this, we set the fraction of clients selected at each round to 0.1, 0.5, and 1.0, respectively. This participation fraction applies to both the expert learning and the global model learning processes. PKD shows consistent improvements across different fractions, increasing the minimum class-wise accuracy by 1.4%, 1.5%, and 1.7%, respectively.
4.3. Comparison with Existing Techniques
@@
Table 6 compares the proposed method with several existing techniques that tackle the ICD problem induced by class imbalance. An effective approach to address the class imbalance issue is to offset the negative impact brought by the uneven class distribution (Cui et al., 2019; Lin et al., 2020). It can be done by either re-sampling the minority classes or assigning larger weights to their loss values while calculating the empirical risk. Here we adopt a similar idea as (Lin et al., 2020) and test its performance for alleviating the inherent ICD problem. Instead of up-weighting the minority classes, we dynamically enhance the weights of the inherent weak classes when calculating the loss value. The result indicates that directly applying this method may not alleviate the inherent ICD problem. It is observed that the minimum class-wise accuracy drops due to increased disparity among weak classes within a group. Take the weak class group {‘shirt’, ‘coat’, ‘pullover’} in Fig. 6 as an example. As mentioned in Section 2.2, an improvement in the accuracy of one weak class is generally accompanied by a corresponding decrease in the accuracy of other weak classes. Suppose at a given epoch, ‘pullover’ is the worst-performing class and assigned a large weight while calculating the total loss. In subsequent iterations, the model is updated to focus more on improving performance for ‘pullover’. However, this improvement may come at the cost of degraded performance on the other two weak classes, thereby resulting in no significant improvement in overall accuracy. In (Devries and Taylor, 2017) and (Shuai et al., 2022), data augmentation is applied at the sample level and feature level, respectively, to implicitly enhance data diversity. These two approaches are commonly used to improve the model’s performance on minority classes. The results suggest that feature-level data augmentation can modestly mitigate the inherent ICD issue, leading to a 2.30% improvement in the accuracy of the worst-performing class.
Table 5. Experimental Results on the FEMNIST Dataset
Dataset FEMNIST Fraction 0.1 0.5 1.0 Method FedAvg PKD $\Delta$ FedAvg PKD $\Delta$ FedAvg PKD $\Delta$ Max. 98.10 98.10 0.00 98.10 98.20 0.10 98.10 98.20 0.10 Ave. 92.30 92.56 0.26 92.78 93.00 0.22 92.96 93.27 0.31 Min. 87.50 89.20 1.70 88.20 89.70 1.50 88.60 90.00 1.40
Table 6. Comparison with Existing ICD Solutions for the Class Imbalance Problem.
Setup FashionMNIST, Local Class Balance, #client=10 Method Fedavg Re-weighting Data Augmentation Ours Focal loss Sample-level Feature-level (McMahan et al., 2017) (Lin et al., 2020) $\Delta$ (Devries and Taylor, 2017) $\Delta$ (Shuai et al., 2022) $\Delta$ PKD $\Delta$ Max. 97.30 97.80 0.50 97.40 0.10 97.40 0.10 97.60 0.30 Ave. 89.29 88.64 -0.65 85.04 -4.25 89.52 0.23 89.50 0.21 Min. 68.40 65.40 -3.00 53.80 -14.60 70.70 2.30 76.60 8.20
Table 7 compares the proposed method with FedProx (Sahu et al., 2018) and FedCodl (Ni et al., 2022) under the Dirichlet partition scenario. These methods are designed to tackle data heterogeneity in FL. Specifically, FedProx introduces a proximal term to the objective to constrain local updates and uses a hyperparameter $\mu$ to control its strength. FedCodl, on the other hand, adopts a KD framework, encouraging local models to align with the global model by minimizing the KL divergence between their output logits. Experimental results show that while existing methods can mitigate general data heterogeneity and improve overall model performance, they exhibit limited effectiveness in enhancing accuracy for weak classes compared to the proposed approach.
Table 7. Comparison with Existing Data-Heterogeneity Solutions
Setup FashionMNIST, Dir(0.5), #client=20 Method Fedavg FedProx(Sahu et al., 2018) FedCodl Ours (McMahan et al., 2017) $\mu=0.001$ $\Delta$ $\mu=0.002$ $\Delta$ (Ni et al., 2022) $\Delta$ PKD $\Delta$ Max. 96.90 97.00 0.10 97.30 0.40 97.50 0.60 96.90 0.00 Ave. 86.05 86.17 0.12 86.06 0.01 86.59 0.54 86.20 0.15 Min. 53.10 56.90 3.80 55.60 2.50 53.80 0.70 63.20 10.10
4.4. Performance on More Complex Datasets
To assess the effectiveness of this approach on more complex datasets with higher class diversity, we further evaluate its performance on CIFAR-100, which comprises 100 classes. We maintain a local class-balanced partition across 10 clients and compare PKD to the baseline. Similar to Fig. 4(c) for FashionMNIST, Fig. 10 maps the predictions to their respective original classes. Given that CIFAR-100 has a large number of classes, the weak class groups are expected to increase in size. For the PKD implementation, we select $G$ groups with the poorest average accuracy. Due to constraints on the number of experts and corresponding overhead, $G$ is set to a small value, specifically 1 or 2, in line with other datasets. The number of weak classes in the two groups is 5 and 4, respectively. The results are shown in Table 8. Despite utilizing only 1-2 expert models, the PKD method achieves a 3-5% improvement in minimum class-wise accuracy.
Table 8. Experimental Results on the CIFAR-100 Dataset.
Setup CIFAR-100, Local Class Balance, #client=10 Method FedAvg PKD $G=1$ $\Delta$ $G=2$ $\Delta$ Max. 91.00 92.00 1.00 92.00 1.00 Ave. 62.82 64.02 1.20 62.91 0.09 Min. 27.00 30.00 3.00 32.00 5.00
Figure 10. Prediction Results of ResNet-56 on CIFAR-100.
4.5. Overhead Analysis
Table 9. Computation Cost in the Local Class-Balanced Scenario (Unit: $U$, which is 374.87 GFLOPs for the Benchmark)
Setup FashionMNIST, Local Class Balance, #client=10 Method Target Min. Acc. Final Acc. 40% 50% 60% 65% 70% Min. Ave. FedAvg (McMahan et al., 2017) 13 19 30 57 - 68.4% 89.3% Focal Loss (Lin et al., 2020) 14 23 56 72 - 65.4% 88.6% Data Aug. (Devries and Taylor, 2017) 29 67 - - - 53.8% 85.0% Data Aug. (Shuai et al., 2022) 14 20 37 63 87 70.7% 89.5% PKD 44.1 50.4 53.5 55.5 76.1 76.6% 89.5%
Table 10. Computation Cost in the Dirichlet Partition Scenario (Unit: Same as Table 9)
Setup FashionMNIST, Dir(0.5), #client=20 Method Target Min. Acc. Final Acc. 30% 45% 50% 55% 60% Min. Ave. FedAvg (McMahan et al., 2017) 16 27 41 - - 53.1% 86.0% FedProx (Sahu et al., 2018) 18 25 36 48 - 56.9% 86.2% FedCodl (Ni et al., 2022) 20 28 35 46 - 55.1% 86.6% PKD 25.2 30.4 35.5 43.6 58.9 63.2% 86.2%
Figure 11. Expert Activation Frequency for FashionMNIST under (a) the Local Class-Balanced Scenario and (b) the Dirichlet Partition $Dir(0.5)$ Scenario.